# ChuLo: Chunk-Level Key Information Representation for Long Document Processing

### v3

@@ 5.4 Token Classification Performance in longer documents
Figures 2(a) and 2(b) visualize the performance breakdown across varying length ranges. For the CoNLL, our model maintains high performance in all length intervals, while Longformer and BigBird exhibit comparable performance within the [1k-2k) range but degrade significantly for longer texts, even for documents that do not exceed their maximum input length. This discrepancy suggests that the uneven distribution of document lengths in their pretraining corpora may lead to inconsistent performance on longer sequences. In contrast, our model’s ability to compress the entire document into 512-length chunks for the decoder enables it to leverage complete contextual information, resulting in better stability and accuracy even on longer documents. For the GUM, where document lengths are shorter (up to 1.3k tokens), our model consistently outperforms Longformer and BigBird in all intervals. The stable performance of all models on GUM aligns with the results on CoNLL, further confirming that our approach’s chunk representation is particularly effective when documents reach lengths that exceed the standard input capacities of the baselines. These results underscore the effectiveness of our chunk representation, which emphasizes keyphrase information, for coarse-grained document classification and fine-grained token-level classification tasks like NER. The ability to maintain performance across varying document lengths highlights the importance of incorporating global contextual information in NER tasks—a largely underexplored aspect. Additionally, off-the-shelf LLMs such as GPT-4o and Gemini 1.5 Pro show suboptimal performance on NER tasks without fine-tuning, and their performance deteriorates further as document length increases. This indicates that, despite their advancements, LLMs still require substantial optimization for token classification tasks and effective long document understanding.
@@ 5.5 Ablation Studies
Keyphrase method HP LUN
Average 0.9538 0.5951
YAKE 0.8769 0.5951
PromptRank 0.9538 0.6440
Sentence Embedding HP LUN
w/o sentence emb. 0.9538 0.6440
sentence emb. 0.9076 0.5537
Backbone HP LUN
BERT (Ours) 0.9538 0.6440
RoBERTa 0.8615 0.5906
Longformer 0.8923 0.5600
Table 8: Effect of different backbone models for the chunk attention.
We performed a few ablation studies on the HP and LUN to assess the impact of various components within our model. First, we analyzed the effectiveness of different keyphrase extraction methods and the effect of using average chunk representations. As shown in Table 8, the PromptRank-based method yields the highest performance across both datasets, outperforming alternatives like YAKE-based. This improvement can be attributed to PromptRank’s ability to extract higher-quality keyphrases by considering semantic relationships within the document, whereas YAKE relies primarily on statistical features such as phrase frequency, resulting in less semantically rich keyphrases. Then, we explored the effect of incorporating sentence embeddings into the chunk representations to introduce global sentence-level context. Surprisingly, as shown in Table 8, the results indicate a performance drop when sentence embeddings are included. We hypothesize that adding sentence-level information at the initial representation stage may cause chunk embeddings within the same sentence to become too similar, hindering the model’s ability to learn distinctive patterns and reducing overall classification performance. We also evaluated the performance of different backbone models for the chunk attention module while keeping the keyphrase extraction and chunk representation settings consistent. Table 8 shows that BERT outperforms Longformer as the backbone. This result suggests that, after document chunking, the input sequences become relatively short, making it difficult for Longformer to leverage its long-range attention capabilities fully. Consequently, Longformer may suffer underutilization during training, resulting in suboptimal performance compared to BERT.
@@ 5.6 Qualitative Analysis
We performed a qualitative analysis by visualizing a sample document from the GUM, comparing the outputs of Longformer, GPT-4o, Gemini 1.5 Pro, and our ChuLo. ChuLo captures the context and semantic patterns of the document, providing accurate predictions, whereas the other models struggle to maintain coherence and consistency. We have more examples in Appendix A.7.
@@ 6 Conclusion
We introduced ChuLo, a novel chunk representation method that enhances the performance of Transformer-based models on long document classification and token-level tasks. By utilizing unsupervised keyphrase extraction, ChuLo effectively reduces input length while preserving critical information, addressing the limitations of truncation and sparse attention. Extensive experiments demonstrate that ChuLo outperforms existing methods by maintaining both global context and high accuracy, even for lengthy inputs. Our results highlight the effectiveness of ChuLo as a robust solution for long document understanding, enabling processing of complex texts in NLP applications.
@@ A.1 Related works
A.2 Baselines
We use BERT (Kenton & Toutanova, 2019) as our backbone model, comparing it with ToBERT (Pappagari et al., 2019), CogLTX (Ding et al., 2020), Longformer (Beltagy et al., 2020), various BERT variants (Park et al., 2022) and ChunkBERT (Jaiswal & Milios, 2023) for the document classification task. For the NER task, we compare against Longformer, BigBird (Zaheer et al., 2020), and two large language models, GPT4o and Gemini1.5pro. Below are brief descriptions of the baseline models:
@@ A.2 Baselines
  • BERT+TextRank and BERT+Random: Proposed to select other tokens randomly or with the help of TextRank(Mihalcea & Tarau, 2004) to feed into the BERT model as the supplementation for long document understanding.
@@ A.3 Baseline Input
BERT (Kenton & Toutanova, 2019) The first 512 tokens
@@
ChunkBERT (Jaiswal & Milios, 2023) The first 4096 tokens
@@ A.4 Details of datasets
A.5 Implementation details
A.5.1 Experiment hyperparameters
@@ A.5.2 Hardware information
A.6 Prompt method
We employed zero-shot prompting with large language models (LLMs), specifically Gemini 1.5 Pro and GPT4o, in our experiments. The prompts used for each dataset are detailed in Table 13 and 14:
Dataset Prompt
LUN Task Definition: You are provided with a news article. Your task is to classify the article into one of the following categories: ”Satire” ”Hoax” or ”Propaganda” Respond only with the appropriate category. The news is: [{input}].
HP Task Definition: You are provided with a news article. Your task is to classify whether the article is hyperpartisan. Respond only with ”True” if the news is hyperpartisan or ”False” if it is not. The news is: [{input}].
Table 13: The prompt we used for each dataset in our experiments.
Dataset Prompt
CoNLL In the task of Named Entity Recognition, the B-, I-, and O- prefixes are commonly used to annotate slot types, indicating the boundaries and types of slots. These labels typically represent: B- (Begin): Signifies the beginning of a slot, marking the start of a new slot. I- (Inside): Represents the interior of a slot, indicating a continuation of the slot. O (Outside): Denotes parts of the input that are not part of any slot. For instance, in a sentence where we want to label a ”date” slot, words containing date information might be tagged as ”B-date” (indicating the beginning of a date slot), followed by consecutive words carrying date information tagged as ”I-date” (indicating the continuation of the date slot), while words not containing date information would be tagged as ”O” (indicating they are outside any slot). Definition: In this task, you are given a conversation, where the words spoken by a person are shown as a list. Your job is to classify the words in the following conversation into one of the 37 different entities. The entities are: ”O”, ”B-PERSON”, ”I-PERSON”, ”B-NORP”, ”I-NORP”, ”B-FAC”, ”I-FAC”, ”B-ORG”, ”I-ORG”, ”B-GPE”, ”I-GPE”, ”B-LOC”, ”I-LOC”, ”B-PRODUCT”, ”I-PRODUCT”, ”B-DATE”, ”I-DATE”, ”B-TIME”, ”I-TIME”, ”B-PERCENT”, ”I-PERCENT”, ”B-MONEY”, ”I-MONEY”, ”B-QUANTITY”, ”I-QUANTITY”, ”B-ORDINAL”, ”I-ORDINAL”, ”B-CARDINAL”, ”I-CARDINAL”, ”B-EVENT”, ”I-EVENT”, ”B-WORK_OF_ART”, ”I-WORK_OF_ART”, ”B-LAW”, ”I-LAW”, ”B-LANGUAGE”, ”I-LANGUAGE”. Only output entities. And the entity types should be output as a list without any explanation. The input is [{input}].
GUM In the task of Named Entity Recognition, the B-, I-, and O- prefixes are commonly used to annotate slot types, indicating the boundaries and types of slots. These labels typically represent: B- (Begin): Signifies the beginning of a slot, marking the start of a new slot. I- (Inside): Represents the interior of a slot, indicating a continuation of the slot. O (Outside): Denotes parts of the input that are not part of any slot. For instance, in a sentence where we want to label a ”date” slot, words containing date information might be tagged as ”B-date” (indicating the beginning of a date slot), followed by consecutive words carrying date information tagged as ”I-date” (indicating the continuation of the date slot), while words not containing date information would be tagged as ”O” (indicating they are outside any slot). Definition: In this task, you are given a conversation, where the words spoken by a person are shown as a list. Your job is to classify the words in the following conversation into one of the 37 different entities. The entities are: ”I-abstract”, ”B-object”, ”B-place”, ”I-substance”, ”I-time”, ”I-place”, ”B-time”, ”B-abstract”, ”I-person”, ”B-plant”, ”B-substance”, ”I-animal”, ”B-organization”, ”I-event”, ”B-person”, ”B-event”, ”I-plant”, ”I-organization”, ”O”, ”I-object”, ”B-animal”. Only output entities. And the entity types should be output as a list without any explanation. The input is [{input}].
Table 14: The prompt we used for each dataset in our experiments.
A.7 More case studies
@@
In this section, we will present several prompt and output samples for the long documents from the LUN (Figures 4) and  5) and Hyperpartisan (Figures 6 and  7) datasets for document classification, as well as GUM (Figures 9 and  10) and CoNLL (Figures 11,  12,  3 and  14) datasets for NER task. Documents with various lengths are randomly selected to see the comparison of our model against GPT-4, Gemini1.5pro and Longformer. While there is always at least one baseline which predicts wrongly for the difficult cases presented for the document classification task, we can observe that our model consistently classifies those documents well. For the token classification task, our model can also correctly classify more tokens than each baseline across the shown cases.

### v4

@@ 5.4 Token Classification in Longer Documents
We further analyze the NER performance across different document length ranges. As presented in Table 5(a) and Table 5(b), we report the number of documents exceeding specific length thresholds and their corresponding performance metrics. On the CoNLL, as document lengths exceed the maximum input capacities of Longformer and BigBird, both models exhibit significant performance drops to 31.56% and 31.45%, respectively. In contrast, our model experiences a minimal decrease of 1.28%, showcasing its resilience and effectiveness in handling long sequences. For the GUM, where all document lengths are within the acceptable range for these models, performance remains stable across all models, with our approach consistently achieving the best results. Figures 2(a) and 2(b) visualize the performance breakdown across varying length ranges. For the CoNLL, our model maintains high performance in all length intervals, while Longformer and BigBird exhibit comparable performance within the [1k-2k) range but degrade significantly for longer texts, even for documents that do not exceed their maximum input length. This discrepancy suggests that the uneven distribution of document lengths in their pretraining corpora may lead to inconsistent performance on longer sequences. In contrast, our model’s ability to compress the entire document into 512-length chunks for the decoder enables it to leverage complete contextual information, resulting in better stability and accuracy even on longer documents. For the GUM, where document lengths are shorter (up to 1.3k tokens), our model consistently outperforms Longformer and BigBird in all intervals. The stable performance of all models on GUM aligns with the results on CoNLL, further confirming that our approach’s chunk representation is particularly effective when documents reach lengths that exceed the standard input capacities of the baselines. These results underscore the effectiveness of our chunk representation, which emphasizes keyphrase information, for coarse-grained document classification and fine-grained token-level classification tasks like NER. The ability to maintain performance across varying document lengths highlights the importance of incorporating global contextual information in NER tasks—a largely underexplored aspect. Additionally, off-the-shelf LLMs such as GPT-4o and Gemini 1.5 Pro show suboptimal performance on NER tasks without fine-tuning, and their performance deteriorates further as document length increases. This indicates that, despite their advancements, LLMs still require substantial optimization for effective long document understanding.
5.5 Prompt Method
We employed zero-shot prompting with large language models (LLMs), specifically Gemini 1.5 Pro and GPT4o, in our experiments. The prompts used for each dataset are detailed in Table 6 and 7:
Dataset Prompt
LUN Task Definition: You are provided with a news article. Your task is to classify the article into one of the following categories: "Satire” "Hoax” or "Propaganda” Respond only with the appropriate category. The news is: [{input}].
HP Task Definition: You are provided with a news article. Your task is to classify whether the article is hyperpartisan. Respond only with "True” if the news is hyperpartisan or "False” if it is not. The news is: [{input}].
Table 6: The prompt we used for each dataset in our experiments.
Dataset Prompt
CoNLL In the task of Named Entity Recognition, the B-, I-, and O- prefixes are commonly used to annotate slot types, indicating the boundaries and types of slots. These labels typically represent: B- (Begin): Signifies the beginning of a slot, marking the start of a new slot. I- (Inside): Represents the interior of a slot, indicating a continuation of the slot. O (Outside): Denotes parts of the input that are not part of any slot. For instance, in a sentence where we want to label a "date" slot, words containing date information might be tagged as "B-date" (indicating the beginning of a date slot), followed by consecutive words carrying date information tagged as "I-date" (indicating the continuation of the date slot), while words not containing date information would be tagged as "O" (indicating they are outside any slot). Definition: In this task, you are given a conversation, where the words spoken by a person are shown as a list. Your job is to classify the words in the following conversation into one of the 37 different entities. The entities are: "O", "B-PERSON", "I-PERSON", "B-NORP", "I-NORP", "B-FAC", "I-FAC", "B-ORG", "I-ORG", "B-GPE", "I-GPE", "B-LOC", "I-LOC", "B-PRODUCT", "I-PRODUCT", "B-DATE", "I-DATE", "B-TIME", "I-TIME", "B-PERCENT", "I-PERCENT", "B-MONEY", "I-MONEY", "B-QUANTITY", "I-QUANTITY", "B-ORDINAL", "I-ORDINAL", "B-CARDINAL", "I-CARDINAL", "B-EVENT", "I-EVENT", "B-WORK_OF_ART", "I-WORK_OF_ART", "B-LAW", "I-LAW", "B-LANGUAGE", "I-LANGUAGE". Only output entities. And the entity types should be output as a list without any explanation. The input is [{input}].
GUM In the task of Named Entity Recognition, the B-, I-, and O- prefixes are commonly used to annotate slot types, indicating the boundaries and types of slots. These labels typically represent: B- (Begin): Signifies the beginning of a slot, marking the start of a new slot. I- (Inside): Represents the interior of a slot, indicating a continuation of the slot. O (Outside): Denotes parts of the input that are not part of any slot. For instance, in a sentence where we want to label a "date" slot, words containing date information might be tagged as "B-date" (indicating the beginning of a date slot), followed by consecutive words carrying date information tagged as "I-date" (indicating the continuation of the date slot), while words not containing date information would be tagged as "O" (indicating they are outside any slot). Definition: In this task, you are given a conversation, where the words spoken by a person are shown as a list. Your job is to classify the words in the following conversation into one of the 37 different entities. The entities are: "I-abstract", "B-object", "B-place", "I-substance", "I-time", "I-place", "B-time", "B-abstract", "I-person", "B-plant", "B-substance", "I-animal", "B-organization", "I-event", "B-person", "B-event", "I-plant", "I-organization", "O", "I-object", "B-animal". Only output entities. And the entity types should be output as a list without any explanation. The input is [{input}].
Table 7: The prompt we used for each dataset in our experiments.
@@ 5.6 Ablation Studies
We conducted more ablation studies, including 1) keyphrase extraction methods, 2) sentence embedding approaches, and 3) backbone model analysis, shown in Appendix A.7.
@@ 5.7 Qualitative Analysis
We performed a qualitative analysis by visualizing each sample document from different datasets, comparing the outputs of Longformer, GPT-4o, Gemini 1.5 Pro, and our ChuLo. ChuLo captures the context and semantic patterns of the document, providing accurate predictions, whereas the other models struggle to maintain coherence and consistency. We have more examples in Appendix A.8.
@@ 6 Conclusion
We introduced ChuLo, a novel chunk representation method that enhances the performance of Transformer-based models on long document classification and token-level classification tasks. By utilizing unsupervised keyphrase extraction, ChuLo effectively reduces input length while preserving critical information, addressing the limitations of truncation and sparse attention. Extensive experiments demonstrate that ChuLo outperforms existing methods by maintaining both global context and high accuracy, even for lengthy inputs. Our research results highlight the effectiveness of ChuLo as a robust solution for long document understanding, enabling processing of complex texts in NLP applications.
7 Limitation
There are several opportunities for future work, including extending the chunk representation to generative tasks such as long text generation, where chunk representation may extend the LLM’s context range limitation and enhance generation quality. However, the performance of the keyphrase extraction method poses a potential risk, as its quality directly affects the overall effectiveness of the approach. We believe this work offers valuable insights into long text understanding and lays a foundation for advancements in related tasks.
Acknowledgments
This study was supported by funding from the Google Award for Inclusion Research Program (G222897).
@@ A.1 Related Works
A.2 Keyphrase Extraction
We employ the Semantic Keyphrase Prioritization (SKP) algorithm to extract keyphrases that encapsulate the key semantic information of the entire document. The detailed are shown in Algorithm 1. While PromptRank uses prompts to rank keyphrases across the first segment of the document determined by its encoder model, our SKP applies this concept at the entire document level to ensure that each chunk can preserve the most informative content for the entire document. After obtaining the sorted phrases set $K_{s}$, we select top-n phrases as the keyphrases of the document, which can be regarded as ranked phrases according to their contextual significance within the entire document.
Algorithm 1 Semantic Keyphrase Prioritization (SKP) Algorithm
[…]
Parameter: Experimentally determined $\alpha$, $\gamma$
Output: Sorted keyphrases set $K_{s}$
1:  Let $S=\emptyset$, $K_{s}=\emptyset$, $i=0$, $j=0$.
2:  Get the candidate phrases set: $K=\mathcal{F}_{REG}(\mathcal{F}_{POS}(D))=\{k_{0},k_{1},\dots,k_{n-1}\}$
3:  Split $D$ into segments $S=\{D_{0},D_{1},\dots,D_{m-1}\}$ to meet the input requirement of $\mathcal{F}_{E}$
4:  for  $i<n$ do
7:     for  $j<m$ do
9:     end for
11:  end for
12:  return $K_{s}=Sort(K,s)$
@@ A.2 Keyphrase Extraction
A.3 Baselines
We use BERT (Kenton and Toutanova, 2019) as our backbone model, comparing it with ToBERT (Pappagari et al., 2019), CogLTX (Ding et al., 2020), Longformer (Beltagy et al., 2020), various BERT variants (Park et al., 2022) and ChunkBERT (Jaiswal and Milios, 2023) for the document classification task. For the NER task, we compare against Longformer, BigBird (Zaheer et al., 2020), and two large language models, GPT4o and Gemini1.5pro. Below are brief descriptions of the baseline models:
@@ A.3 Baselines
  • BERT+TextRank and BERT+Random: Proposed to select other tokens randomly or with the help of TextRank(Mihalcea and Tarau, 2004) to feed into the BERT model as the supplementation for long document classification.
@@ A.4 Baseline Input
BERT (Kenton and Toutanova, 2019) The first 512 tokens
@@
ChunkBERT (Jaiswal and Milios, 2023) The first 4096 tokens
@@ A.5 Details of datasets
A.6 Implementation details
A.6.1 Experiment hyperparameters
@@ A.6.2 Hardware Information
A.7 Ablation Studies
We performed a few ablation studies on the HP and LUN to assess the impact of various components within our model. First, we analyzed the effectiveness of different keyphrase extraction methods and the effect of using average chunk representations. As shown in Table 12, the PromptRank-based method yields the highest performance across both datasets, outperforming alternatives like YAKE-based. This improvement can be attributed to PromptRank’s ability to extract higher-quality keyphrases by considering semantic relationships within the document, whereas YAKE relies primarily on statistical features such as phrase frequency, resulting in less semantically rich keyphrases. Then, we explored the effect of incorporating sentence embeddings into the chunk representations to introduce global sentence-level context. Surprisingly, as shown in Table 13, the results indicate a performance drop when sentence embeddings are included. We hypothesize that adding sentence-level information at the initial representation stage may cause chunk embeddings within the same sentence to become too similar, hindering the model’s ability to learn distinctive patterns and reducing overall classification performance. Then, we evaluated the performance of different backbone models for the chunk attention module while keeping the keyphrase extraction and chunk representation settings consistent. Table 14 shows that BERT outperforms Longformer as the backbone. This result suggests that, after document chunking, the input sequences become relatively short, making it difficult for Longformer to leverage its long-range attention capabilities fully. Consequently, Longformer may suffer underutilization during training, resulting in suboptimal performance compared to BERT.
       Keyphrase method        HP        LUN
       Average        0.9538        0.5951
       YAKE        0.8769        0.5951
       PromptRank        0.9538        0.6440
       Sentence Embedding        HP        LUN
       w/o sentence emb.        0.9538        0.6440
       sentence emb.        0.9076        0.5537
       Backbone        HP        LUN
       BERT (Ours)        0.9538        0.6440
       RoBERTa        0.8615        0.5906
       Longformer        0.8923        0.5600
Table 14: Effect of different backbone models for the chunk attention.
A.8 More Case Studies
@@
In this section, we will present several prompt and output samples for the long documents from the LUN (Figures 4) and  5) and Hyperpartisan (Figures 6 and  7) datasets for document classification, as well as GUM (Figures 8, 9 and  10) and CoNLL (Figures 11,  3,  12 and  13) datasets for NER task. Documents with various lengths are randomly selected to see the comparison of our model against GPT-4, Gemini1.5pro and Longformer. While there is always at least one baseline which predicts wrongly for the difficult cases presented for the document classification task, we can observe that our model consistently classifies those documents well. For the token classification task, our model can also correctly classify more tokens than each baseline across the shown cases.